Consider a firm with a production function given by Q = 40N − 0.125N^2 . This implies that the marginal product of labour is MPN = 40 − 0.25N. Suppose that the market price for the good that the firm produces is P = 1, and the market wage is w = 5.

(a) How many workers will the firm choose to hire?

b) Compute the firm’s profits given the optimal choice of employment

c) At what wage level would the firm be willing to hire exactly 100 workers?
Answer #1

a. At optimum MPN= w/P

P= 1

40 - 0.25N = 5

N= 140

Number of workers hired is 140

b. Q= 40*140 - 0.125*(140)^2= 3150

Profit = 3150 - 5*140= 2450

c. N= 100

40-0.25*100= 15